OpenStudy (anonymous):

Eliminate the parameter. x = 4 cos t, y = 4 sin t

OpenStudy (amistre64):

solve for t in x, then replace t by t(x) in y

OpenStudy (anonymous):

Alternatively, you can take the fact that the parametric equations describe a circle with radius 4, so you have \[(4\cos t)^2+(4\sin t)^2=4^2\\ x^2+y^2=16\]